# YOLOv8-AM: YOLOv8 with Attention Mechanisms for Pediatric Wrist Fracture Detection

### v4

@@ Abstract
Keywords:
Deep Learning Fracture Detection Object Detection Medical Image Processing Radiology X-ray Imaging.
1 Introduction
@@
Computer-assisted diagnosis (CAD) of medical images provides experts (radiologists, surgeons, etc.) with help in some decision tasks. With the continuous development of deep learning and the improvement of medical image processing techniques [1, 9, 10, 53], more and more researchers are trying to employ neural networks for CAD, including fracture detection [3, 15, 27, 37, 65].
You Only Look Once (YOLO) model [46], as one of the most important models for object detection tasks [26], demonstrates satisfactory performance in fracture detection [19, 51, 52]. With the introduction of YOLOv8 [23], the latest version of the YOLO models, by Ultralytics in 2023, it has been widely used in various object detection tasks. As shown in Fig. 1, GRAZPEDWRI-DX [43] is a public dataset of 20,327 pediatric wrist trauma X-ray images. This work follows the study [25] that evaluates the performance of different network models on this dataset in a fracture detection task.
@@
Therefore, we propose the YOLOv8-AM model for fracture detection by employing four different attention modules, including Convolutional Block Attention Module (CBAM) [63], Global Attention Mechanism (GAM) [39], Efficient Channel Attention (ECA) [59] and Shuffle Attention (SA) [69], to the YOLOv8 architecture, as shown in Fig. 2.
Figure 1: Examples of the input images and ground truth images used in this study.
@@
This paper is organized as follows: Section 2 introduces the research on fracture detection utilizing deep learning methods and outlines the evolution of attention mechanism. Section 3 presents the overall architecture of the YOLOv8-AM model and four different attention modules employed. Section 4 conducts a comparative analysis of the performance of four YOLOv8-AM models against the baseline YOLOv8 model. Subsequently, Section 5 explores the reasons why GAM provides less improvement in the performance of the YOLOv8-AM model for fracture detection. Finally, Section 6 concludes this research work.
2 Related Work
2.1 Fracture Detection
Fracture detection is a hot topic in medical image processing (MIP). Researchers usually employ various neural networks for prediction, including the YOLO series models [23, 24, 46]. Burkow et al. [6] utilized the YOLOv5 model [24] to recognize rib fracture on 704 pediatric Chest X-ray (CXR) images. Tsai et al. [54] performed data augmentation on CXR images and subsequently employed the YOLOv5 model to detect fractures. Warin et al. [61] firstly categorized the maxillofacial fractures into four categories (frontal, midfacial, mandibular, and no fracture), and predicted them using the YOLOv5 model on 3,407 CT images. Additionally, Warin et al. [60] used the YOLOv5 model to detect fractures in the X-ray images of the mandible. Yuan et al. [66] incorporated external attention and 3D feature fusion methods into the YOLOv5 model for fracture detection in skull CT images. Furthermore, vertebral localization proves valuable for recognizing vertebral deformities and fractures. Mushtaq et al. [42] utilized YOLOv5 for lumbar vertebrae localization, achieving the mean Average Precision (mAP) value of 0.957. Although the YOLOv5 model is extensively employed in fracture detection, the utilization of the YOLOv8 model [23] is comparatively rare.
2.2 Attention Module
@@
Figure 3: Detailed illustration of the YOLOv8-AM model architecture, where the attention module is Shuffle Attention (SA), Efficient Channel Attention (ECA), Global Attention Mechanism (GAM), and ResBlock + Convolutional Block Attention Module (ResCBAM).
3 Methodology
3.1 Baseline Model
YOLOv8 architecture comprises four key components: Backbone, Neck, Head, and Loss Function. The Backbone incorporates the Cross Stage Partial (CSP) [57] concept, offering the advantage of reducing computational loads while enhancing the learning capability of CNNs. Illustrated in Fig. 3, YOLOv8 diverges from YOLOv5 employing the C3 module [24], adopting the C2f module, which integrates the C3 module and the Extended ELAN [58] (E-ELAN) concept from YOLOv7 [56]. Specifically, the C3 module involves three convolutional modules and multiple bottlenecks, whereas the C2f module consists of two convolutional modules concatenated with multiple bottlenecks. The convolutional module is structured as Convolution-Batch Normalization-SiLU (CBS).
In the Neck part, YOLOv5 employs the Feature Pyramid Network (FPN) [36] architecture for top-down sampling, ensuring that the lower feature map incorporates richer feature information. Simultaneously, the Path Aggregation Network (PAN) [38] structure is applied for bottom-up sampling, enhancing the top feature map with more precise location information. The combination of these two structures is executed to guarantee the accurate prediction of images across varying dimensions. YOLOv8 follows the FPN and PAN frameworks while deleting the convolution operation during the up-sampling stage, as illustrated in Fig. 3.
@@ 3.1 Baseline Model
$Loss_{DF}=-[(y_{n+1}-y)\log{\frac{y_{n+1}-y_{n}}{y_{n+1}-y_{n}}}+(y-y_{n})\log{\frac{y-y_{n}}{y_{n+1}-y_{n}}}].$ (2)
The CIoU Loss introduces an influence factor to the Distance Intersection over Union (DIoU) Loss [72] by considering the aspect ratio of the predicted bounding box and the ground truth bounding box. The corresponding equation is as follows:
@@
where $IoU$ measures the overlap between the predicted and ground truth bounding boxes; $d$ is the Euclidean distance between the center points of the predicted and ground truth bounding boxes, and $c$ is the diagonal length of the smallest enclosing box that contains both predicted and ground truth bounding boxes. Additionally, $v$ represents the parameter quantifying the consistency of the aspect ratio, defined by the following equation:
@@
where $w$ denotes the weight of the bounding box; $h$ represents the height of the bounding box; $gt$ means the ground truth, and $p$ means the prediction.
@@ 3.2 Proposed Method
This work incorporates the attention module into the Neck part of YOLOv8 to enhance the capture of key features and suppress the interfering information. As illustrated in Fig. 3, attention modules, namely CBAM [63], GAM [39], ECA [59], and SA [69], are independently employed after each of the four C2f modules. A detailed introduction of these four attention modules is presented in Section 3.3.
3.3 Attention Modules
3.3.1 Convolutional Block Attention Module (CBAM)
CBAM comprises both channel attention (C-Attention) and spatial attention (S-Attention), as shown on the left of Fig. 4. Given an intermediate feature map denoted as $F_{input}\in\mathbb{R}_{C\texttimes H\texttimes W}$, CBAM sequentially infers a 1D channel attention map $M_{C}\in\mathbb{R}^{C\times 1\times 1}$ and a 2D spatial attention map $M_{S}\in\mathbb{R}^{1\times H\times W}$ through the following equation:
@@ 3.3.1 Convolutional Block Attention Module (CBAM)
It can be seen from the right of Fig. 4, for ResBlock + CBAM (ResCBAM), $F_{output}$ is the element-wise summation of $F_{input}$ and $F_{FR}$ as shown in the following equation:
@@
3.3.2 Efficient Channel Attention (ECA)
ECA primarily encompasses cross-channel interaction and 1D convolution with an adaptive convolution kernel, as shown on the right of Fig. 5. Cross-channel interaction represents a novel approach to combining features, enhancing the expression of features for specific semantics. The input feature map $F_{input}\in\mathbb{R}^{C\times H\times W}$ obtains the aggregated feature $F_{a}$ after performing GAP and cross-channel interaction, where $\mathbb{C}$ represents cross-channel interaction. The equation is shown as follows:
@@ 3.3.2 Efficient Channel Attention (ECA)
For the aggregated features, ECA captures the local cross-channel interaction by considering the interaction between the features of each channel and their neighboring $k$ channels, avoiding the use of 1D convolution for dimensionality reduction, and effectively realizing the multi-channel interaction, where the weights of the features $F_{ai}$ can be calculated as:
@@
3.3.3 Shuffle Attention (SA)
SA divides the input feature maps into different groups, employing the Shuffle Unit to integrate both channel attention and spatial attention into one block for each group, as shown on the left of Fig. 5. Subsequently, the sub-features are aggregated, and the “Channel Shuffle“ operator, as employed in ShuffleNetV2 [40], is applied to facilitate information communication among various sub-features.
@@ 3.3.3 Shuffle Attention (SA)
It can be seen on the left of Fig. 5, the output sub-feature ${x_{k}}^{\prime}$ is obtained by concatenating ${x_{k1}}^{\prime}$ and ${x_{k2}}^{\prime}$, and the equation is shown as follows:
@@
3.3.4 Global Attention Mechanism (GAM)
GAM adopts the main architecture proposed by CBAM consisting of channel at-tention and spatial attention, and redesigns the submodules as shown in Fig. 6. Additionally, we add Shortcut Connection [17] between the layers within GAM, which allows the inputs to propagate forward faster, as shown in the following equation:
$F_{output}=F_{input}+[M_{S}(M_{C}(F_{input})\otimes F_{input})\otimes(M_{C}(F_{input})\otimes F_{input})].$ (18)
@@ 3.3.4 Global Attention Mechanism (GAM)
$M_{C}(F)=\sigma[ReversePermutation(MLP(Permutation(F)))].$ (19)
@@
4 Experiment
4.1 Dataset
4.2 Preprocessing and Data Augmentation
4.3 Evaluation Metric
4.3.1 Parameters (Params)
4.3.2 Floating Point Operations (FLOPs)
Floating Point Operations per Second serves as a metric for assessing the performance of computers or computing systems and is commonly employed to evaluate the computational complexity of neural network models. FLOPs represent the number of floating-point operations executed per second, providing a crucial indicator of the computational performance and speed of the model. In resource-limited environments, models with lower FLOPs may be more suitable, while those with higher FLOPs may necessitate more powerful hardware support.
4.3.3 Mean Average Precision (mAP)
4.3.4 F1 Score
The F1 score is based on the summed average of the precision and recall values of the model. Its value ranges from 0 to 1, where value closer to 1 indicating that the model has a better balance between precision and recall. When one of precision and recall values is biased towards 0, the F1 score will also be close to 0, which indicates poor model performance. Considering both precision and recall values together, the F1 score helps to assess the accuracy of the model in positive category prediction and its sensitivity to positive categories.
In the problem involving imbalanced category classification, relying solely on accuracy for evaluation may lead to bias, as the model may perform better in predicting the category with a larger number of samples. Therefore, using the F1 score provides a more comprehensive assessment of the model performance, especially when dealing with imbalanced datasets.
4.4 Experiment Setup
Table 1: Experimental results of fracture detection on the GRAZPEDWRI-DX dataset using the YOLOv8-AM models with three different attention modules, where Inference means the total time of prediction per X-ray image, and the total time includes the preprocessing, inference, and post-processing time.
Model
Size
Input
Size
mAPval
50
mAPval
50-95
Params
FLOPs
Inference
@@ 4.4 Experiment Setup
Model
Size
Input
Size
mAPval
50
mAPval
50-95
Params
FLOPs
Inference
@@
Model
Size
Input
Size
mAPval
50
mAPval
50-95
Params
FLOPs
Inference
@@
Table 2: Experimental results of fracture detection on the GRAZPEDWRI-DX dataset using the YOLOv8-AM models with GAM and ResGAM, where Inference means the total time of prediction per X-ray image, and the total time includes the preprocessing, inference, and post-processing time.
Model
Size
Input
Size
mAPval
50
mAPval
50-95
Params
FLOPs
Inference
@@
Model
Size
Input
Size
mAPval
50
mAPval
50-95
Params
FLOPs
Inference
@@
Table 3: Quantitative comparison (F1 Score/mAP) of fracture detection on the GRAZPEDWRI-DX dataset using YOLOv8 and YOLOv8-AM models, where Inference means the total time of prediction per X-ray image, and the total time includes the preprocessing, inference, and post-processing time.
@@
Figure 7: Examples of different YOLOv8-AM models for pediatric wrist fracture detection, (a) manually labeled data, (b) ResCBAM, (c) ECA, (d) SA, (e) GAM, (f) ResGAM.
Figure 8: Detailed illustrations of the Precision-Recall Curves (%) for four different YOLOv8-AM models on each category in the GRAZPEDWRI-DX dataset.
4.5 Experimental Results
@@
As shown in Tables 1(c) and 2(b), the performance of the models trained using the training set with input image size of 1024 surpasses that of models trained using the training set with input image size of 640. Nevertheless, it is noteworthy that this improvement in performance is accompanied by an increase in inference time. For example, considering the ResCBAM-based YOLOv8-AM model with the large model size, the mean Average Precision at IoU 50 (mAP 50) attains 42.2% for the input image size of 1024, which is 5.24% higher than that of 40.1% obtained for the input image size of 640. However, the inference time of the model increases from 4.1ms to 8.7ms because the model size becomes larger.
@@
Table 1(c) shows the model performance of the YOLOv8-AM, employing three different attention modules including ResCBAM, SA and ECA. These experimental results are obtained with different model sizes and different input image sizes. In Table 2(b), the model performance of the YOLOv8-AM is presented when utilizing GAM directly. Additionally, we introduce a novel approach by incorporating ResBlock and GAM, named ResGAM, to enhance the overall model performance of the YOLOv8-AM. Specifically, when the input image size is 1024 and the model size is medium, the newly introduced ResGAM demonstrates a notable enhancement in mAP for the YOLOv8-AM model based on GAM. The mAP 50 increases from 63.7% to 64.9%, providing that our proposed ResGAM positively contributes to the performance enhancement.
@@
To compare the effect of different attention modules on the model performance, we organize the experimental data with input image size of 1024 and model size of large. The corresponding results are presented in Table 3. In summary, the F1 score, mAP 50-95, and mAP 50 values for all YOLOv8-AM models surpass that of the YOLOv8 model. Specifically, the mAP 50 for the YOLOv8-AM models based on SA and ECA stands at 64.3% and 64.2%, respectively. These values are marginally superior to 63.6% mAP 50 obtained by the YOLOv8 model. Notably, this enhanced performance requires almost the same inference time as that of the YOLOv8 model. For the YOLOv8-AM model based on ResCBAM, it obtained mAP 50 of 65.8%, achieving the SOTA model performance. However, the gain on the performance of the YOLOv8 model by GAM is not satisfactory, so we propose the incorporation of ResGAM into the YOLOv8-AM model.
@@
In this paper, to evaluate the gain of the attention module on the accuracy of the YOLOv8 model for predicting fractures in a real-world diagnostic scenario, four X-ray images are randomly selected, and Fig. 7 demonstrates the prediction results of different YOLOv8-AM models. The YOLOv8-AM model, serving as a CAD tool, plays a crucial role in supporting radiologists and surgeons during diagnosis by effectively identifying fractures and detecting metal punctures in singular fracture scenarios. However, it is important to note that the accuracy of model prediction may decrease in instances involving dense fractures.
@@
Fig. 8 shows the Precision-Recall Curve (PRC) for each category predicted by different YOLOv8-AM models. From the figure, we can see that different YOLOv8-AM models have greater ability to correctly detect fracture, metal, and text, with the average accuracy exceeding 90%. However, for two categories, bone anomaly and soft tissue, the ability to correctly detect them is poorer, with accuracy approximately at 45% and 30%, respectively. These low accuracy rates seriously effect the mAP 50 values of the models. We consider this is due to the small number of objects within these two categories in the used dataset. As described in GRAZPEDWRI-DX [43], the number of bone anomaly and soft tissue accounts for 0.41% and 0.68% of the total number of objects, respectively. Consequently, any improvement in model performance via architectural enhancements is constrained by this data limitation. To enhance the performance of the model, a recourse to incorporating extra data becomes imperative.
5 Discussion

### v5

@@ Abstract
Index Terms:
Computer Vision, Deep Learning, Fracture Detection, Medical Image Diagnostics, Medical Image Processing, Object Detection, Radiology, X-ray Imaging, You Only Look Once (YOLO).
[…]
=-15pt
I Introduction
@@
Figure 1: Example diagrams illustrating the pediatric wrist fracture prediction task conducted in this work, with the input image displayed on the left of each group and the prediction result shown on the right of each group.
@@
Computer-assisted diagnosis (CAD) of medical images provides experts (i.e., radiologists, surgeons, etc.) with help in some decision tasks. With the continuous development of deep learning and the improvement of medical image processing techniques [12, 13, 14, 15], more and more researchers are trying to employ neural networks for CAD, including fracture detection [16, 17, 18, 19, 20].
You Only Look Once (YOLO) model [21], as one of the most important models for object detection tasks [22], demonstrates satisfactory performance in fracture detection [23, 24, 25]. With the introduction of YOLOv8 [26], the latest version of the YOLO models, by Ultralytics in 2023, it has been widely used in various object detection tasks. As shown in FIGURE 1, GRAZPEDWRI-DX [27] is a public dataset of 20,327 pediatric wrist trauma X-ray images. This work follows the study [28] that evaluates the performance of different network models on this dataset in a fracture detection task.
@@
Therefore, we propose the YOLOv8-AM model for fracture detection by employing four different attention modules, including Convolutional Block Attention Module (CBAM) [38], Global Attention Mechanism (GAM) [39], Efficient Channel Attention (ECA) [40] and Shuffle Attention (SA) [41], to the YOLOv8 architecture, as shown in FIGURE 2.
@@
This paper is organized as follows: Section II introduces the research on fracture detection utilizing deep learning methods and outlines the evolution of attention mechanism. Section III presents the overall architecture of the YOLOv8-AM model and four different attention modules employed. Section IV conducts a comparative analysis of the performance of four YOLOv8-AM models against the baseline YOLOv8 model. Subsequently, Section V discusses the reasons why GAM provides less improvement in the performance of the YOLOv8-AM model for fracture detection. Finally, Section VI concludes this research work and explores the future works.
II Related Work
II-A Fracture Detection
Fracture detection is a hot topic in medical image processing (MIP). Researchers usually employ various neural networks for prediction, including the YOLO series models [26, 43, 21]. Burkow et al. [44] utilized the YOLOv5 model [43] to recognize rib fracture on 704 pediatric Chest X-ray (CXR) images. Tsai et al. [45] performed data augmentation on CXR images and subsequently employed the YOLOv5 model to detect fractures. Warin et al. [46] firstly categorized the maxillofacial fractures into four categories (i.e., frontal, midfacial, mandibular, and no fracture), and predicted them using the YOLOv5 model on 3,407 CT images. In addition, Warin et al. [47] used the YOLOv5 model to detect fractures in the X-ray images of the mandible. Yuan et al. [48] incorporated external attention and 3D feature fusion methods into the YOLOv5 model for fracture detection in skull CT images. Furthermore, vertebral localization proves valuable for recognizing vertebral deformities and fractures. Mushtaq et al. [49] utilized YOLOv5 for lumbar vertebrae localization, achieving the mean Average Precision (mAP) value of 0.957. Although the YOLOv5 model is extensively employed in fracture detection, the utilization of the YOLOv8 model [26] is comparatively rare.
II-B Attention Module
@@
Figure 3: Detailed illustration of the YOLOv8-AM model architecture, where the attention modules are Shuffle Attention (SA), Efficient Channel Attention (ECA), Global Attention Mechanism (GAM), and ResBlock + Convolutional Block Attention Module (ResCBAM), respectively.
III Methodology
III-A Baseline Model
YOLOv8 architecture comprises four key components: Backbone, Neck, Head, and Loss Function. The Backbone incorporates the Cross Stage Partial (CSP) [53] concept, offering the advantage of reducing computational loads while enhancing the learning capability of CNNs. Illustrated in FIGURE 3, YOLOv8 diverges from YOLOv5 employing the C3 module [43], adopting the C2f module, which integrates the C3 module and the Extended ELAN [54] (E-ELAN) concept from YOLOv7 [55]. Specifically, the C3 module involves three convolutional modules and multiple bottlenecks, whereas the C2f module consists of two convolutional modules concatenated with multiple bottlenecks. The convolutional module is structured as Convolution-Batch Normalization-SiLU (CBS).
In the Neck part, YOLOv5 employs the Feature Pyramid Network (FPN) [56] architecture for top-down sampling, ensuring that the lower feature map incorporates richer feature information. Simultaneously, the Path Aggregation Network (PAN) [57] structure is applied for bottom-up sampling, enhancing the top feature map with more precise location information. The combination of these two structures is executed to guarantee the accurate prediction of images across varying dimensions. YOLOv8 follows the FPN and PAN frameworks while deleting the convolution operation during the up-sampling stage, as illustrated in FIGURE 3.
@@ III-A Baseline Model
$\begin{split}Loss_{DF}=&-[(y_{n+1}-y)\log{\frac{y_{n+1}-y_{n}}{y_{n+1}-y_{n}}}\\ &+(y-y_{n})\log{\frac{y-y_{n}}{y_{n+1}-y_{n}}}].\end{split}$ (2)
The CIoU Loss introduces an influence factor to the Distance Intersection over Union (DIoU) Loss [62] by considering the aspect ratio of the predicted bounding box and the Ground-Truth bounding box. The corresponding equation is as follows:
@@
where $IoU$ measures the overlap between the predicted and Ground-Truth bounding boxes; $d$ is the Euclidean distance between the center points of the predicted and Ground-Truth bounding boxes, and $c$ is the diagonal length of the smallest enclosing box that contains both predicted and Ground-Truth bounding boxes. Additionally, $v$ represents the parameter quantifying the consistency of the aspect ratio, defined by the following equation:
@@
where $w$ denotes the weight of the bounding box; $h$ represents the height of the bounding box; $gt$ means the Ground-Truth, and $p$ means the prediction.
@@ III-B Proposed Method
This work incorporates the attention module into the Neck part of YOLOv8 to enhance the capture of key features and suppress the interfering information. As illustrated in FIGURE 3, attention modules, namely CBAM [38], GAM [39], ECA [40], and SA [41], are independently employed after each of the four C2f modules. A detailed introduction of these four attention modules is presented in Section III-C.
III-C Attention Modules
III-C1 Convolutional Block Attention Module (CBAM)
CBAM comprises both channel attention (C-Attention) and spatial attention (S-Attention), as shown on the left of FIGURE 4. Given an intermediate feature map denoted as $F_{input}\in\mathbb{R}_{C\times H\times W}$, CBAM sequentially infers a 1D channel attention map $M_{C}\in\mathbb{R}^{C\times 1\times 1}$ and a 2D spatial attention map $M_{S}\in\mathbb{R}^{1\times H\times W}$ through the following equation:
@@ III-C1 Convolutional Block Attention Module (CBAM)
It can be seen from the right of FIGURE 4, for ResBlock + CBAM (ResCBAM), $F_{output}$ is the element-wise summation of $F_{input}$ and $F_{FR}$ as shown in the following equation:
@@
III-C2 Efficient Channel Attention (ECA)
ECA primarily encompasses cross-channel interaction and 1D convolution with an adaptive convolution kernel, as shown on the right of FIGURE 5. Cross-channel interaction represents a novel approach to combining features, enhancing the expression of features for specific semantics. The input feature map $F_{input}\in\mathbb{R}^{C\times H\times W}$ obtains the aggregated feature $F_{a}$ after performing GAP and cross-channel interaction, where $\mathbb{C}$ represents cross-channel interaction. The equation is shown as follows:
@@ III-C2 Efficient Channel Attention (ECA)
For the aggregated features, ECA captures the local cross-channel interaction by considering the interaction between the features of each channel and their neighboring $k$ channels, avoiding the use of 1D convolution for dimensionality reduction, and effectively realizing the multi-channel interaction, where the weights of features $F_{ai}$ are as follows:
@@
III-C3 Shuffle Attention (SA)
SA divides the input feature maps into different groups, employing the Shuffle Unit to integrate both channel attention and spatial attention into one block for each group, as shown on the left of FIGURE 5. Subsequently, the sub-features are aggregated, and the “Channel Shuffle“ operator, as employed in ShuffleNetV2 [69], is applied to facilitate information communication among various sub-features.
@@ III-C3 Shuffle Attention (SA)
It can be seen on the left of FIGURE 5, the output sub-feature ${x_{k}}^{\prime}$ is obtained by concatenating ${x_{k1}}^{\prime}$ and ${x_{k2}}^{\prime}$, and the equation is shown as follows:
@@
III-C4 Global Attention Mechanism (GAM)
GAM adopts the main architecture proposed by CBAM consisting of channel at-tention and spatial attention, and redesigns the submodules as shown in FIGURE 6. Additionally, we add Shortcut Connection [42] between the layers within GAM, which allows the inputs to propagate forward faster, as shown in the following equation:
$\begin{split}F_{output}=F_{input}+&[M_{S}(M_{C}(F_{input})\otimes F_{input})\\ &\otimes(M_{C}(F_{input})\otimes F_{input})].\end{split}$ (18)
@@ III-C4 Global Attention Mechanism (GAM)
$\begin{split}M_{C}(F)=\sigma&[ReversePermutation(\\ &MLP(Permutation(F)))].\end{split}$ (19)
@@
TABLE I: Experimental results of fracture detection on the GRAZPEDWRI-DX dataset using the YOLOv8-AM models with three different attention modules (i.e., ResCBAM, SA, and ECA).
[…]
@@
[…]
@@
[…]
@@
TABLE II: Experimental results of fracture detection on the GRAZPEDWRI-DX dataset using the YOLOv8-AM models with two attention modules (i.e., GAM and ResGAM).
[…]
@@
[…]
@@
TABLE III: Quantitative comparison (F1-Score/mAP/Inference) of fracture detection on the GRAZPEDWRI-DX dataset using YOLOv8 and YOLOv8-AM models. Best and 2nd best performance are in red and blue colors, respectively.
@@
Figure 7: Examples of results of different YOLOv8-AM models applied to pediatric wrist fracture detection and Ground-Truth, where (a) manually labeled data; (b) ResCBAM; (c) ECA; (d) SA; (e) GAM; and (f) ResGAM.
Figure 8: Detailed illustrations of the Precision-Recall Curves (%) for four different YOLOv8-AM models on each category in the GRAZPEDWRI-DX dataset, where (a) SA; (b) ECA; (c) ResCBAM; and (d) ResGAM.
IV Experiment
IV-A Dataset
IV-B Preprocessing and Data Augmentation
IV-C Evaluation Metric
IV-C1 Parameters (Params)
IV-C2 Floating Point Operations (FLOPs)
Floating Point Operations serves as a metric for assessing the performance of computers or computing systems and is commonly employed to evaluate the computational complexity of neural network models. FLOPs represent the number of floating-point operations executed per second, providing a crucial indicator of the computational performance and speed of the model. In resource-limited environments, models with lower FLOPs may be more suitable, while those with higher FLOPs may necessitate more powerful hardware support.
IV-C3 Mean Average Precision (mAP)
IV-C4 F1-Score
The F1-Score is based on the summed average of the precision and recall values of the model. Its value ranges from 0 to 1, where value closer to 1 indicating that the model has a better balance between precision and recall. When one of precision and recall values is biased towards 0, the F1-Score will also be close to 0, which indicates poor model performance. Considering both precision and recall values together, the F1-Score helps to assess the accuracy of the model in positive category prediction and its sensitivity to positive categories.
In the problem involving imbalanced category classification, relying solely on accuracy for evaluation may lead to bias, as the model may perform better in predicting the category with a larger number of samples. Therefore, using the F1-Score provides a more comprehensive assessment of the model performance, especially when dealing with imbalanced datasets.
IV-C5 Inference Time
Inference Time refers to the duration required by network models to process X-ray images from input to the final prediction, including preprocessing, inference, and post-processing stages. In this work, the inference time per image is measured using single NVIDIA GeForce RTX 3090 GPU.
IV-D Experiment Setup
IV-E Experimental Results
@@
As shown in TABLEs I(c) and II(b), the performance of the models trained using the training set with input image size of 1024 surpasses that of models trained using the training set with input image size of 640. Nevertheless, it is noteworthy that this improvement in performance is accompanied by an increase in inference time. For example, considering the ResCBAM-based YOLOv8-AM model with the large model size, the mean Average Precision at IoU 50 (mAP 50) attains 42.2% for the input image size of 1024, which is 5.24% higher than that of 40.1% obtained for the input image size of 640. However, the inference time of the model increases from 4.1ms to 8.7ms because the model size becomes larger.
@@
TABLE I(c) shows the model performance of the YOLOv8-AM, employing three different attention modules including ResCBAM, SA and ECA. These experimental results are obtained with different model sizes and different input image sizes. In TABLE II(b), the model performance of the YOLOv8-AM is presented when utilizing GAM directly. Additionally, we introduce a novel approach by incorporating ResBlock and GAM, named ResGAM, to enhance the overall model performance of the YOLOv8-AM. Specifically, when the input image size is 1024 and the model size is medium, the newly introduced ResGAM demonstrates a notable enhancement in mAP for the YOLOv8-AM model based on GAM. The mAP 50 increases from 63.7% to 64.9%, providing that our proposed ResGAM positively contributes to the performance enhancement.
@@
To compare the effect of different attention modules on the model performance, we organize the experimental data with input image size of 1024 and model size of large. The corresponding results are presented in TABLE III. In summary, the F1-Score, mAP 50-95, and mAP 50 values for all YOLOv8-AM models surpass that of the YOLOv8 model. Specifically, the mAP 50 for the YOLOv8-AM models based on SA and ECA stands at 64.3% and 64.2%, respectively. These values are marginally superior to 63.6% mAP 50 obtained by the YOLOv8 model. Notably, this enhanced performance requires almost the same inference time as that of the YOLOv8 model. For the YOLOv8-AM model based on ResCBAM, it obtained mAP 50 of 65.8%, achieving the SOTA model performance. However, the gain on the performance of the YOLOv8 model by GAM is not satisfactory, so we propose the incorporation of ResGAM into the YOLOv8-AM model.
@@
In this paper, to evaluate the gain of the attention module on the accuracy of the YOLOv8 model for predicting fractures in a real-world diagnostic scenario, four X-ray images are randomly selected, and FIGURE 7 demonstrates the prediction results of different YOLOv8-AM models. The YOLOv8-AM model, serving as a CAD tool, plays a crucial role in supporting radiologists and surgeons during diagnosis by effectively identifying fractures and detecting metal punctures in singular fracture scenarios. However, it is important to note that the accuracy of model prediction may decrease in instances involving dense fractures.
@@
FIGURE 8 shows the Precision-Recall Curve (PRC) for each category predicted by different YOLOv8-AM models. From the figure, we can see that different YOLOv8-AM models have greater ability to correctly detect fracture, metal, and text, with the average accuracy exceeding 90%. However, for two categories, bone anomaly and soft tissue, the ability to correctly detect them is poorer, with accuracy approximately at 45% and 30%, respectively. These low accuracy rates seriously effect the mAP 50 values of the models. We consider this is due to the small number of objects within these two categories in the used dataset. As described in GRAZPEDWRI-DX [27], the number of bone anomaly and soft tissue accounts for 0.41% and 0.68% of the total number of objects, respectively. Consequently, any improvement in model performance via architectural enhancements is constrained by this data limitation. To enhance the performance of the model, a recourse to incorporating extra data becomes imperative.
V Discussion